# Xinyuan Lu1    Liangming Pan 2    Yubo Ma3    Preslav Nakov4    Min-Yen Kan1

### v2

@@ 5. SCITAB Table Fact Checking Scientific Articles Table, Claim Label (Short)
Table 4: Statistics about the seed datasets for TART training, highlighting their respective tasks, domains, and the nature of input and output data.
@@
Table 5: Statistics about dataset ToolTab for training the TART model.
@@ Appendix B Different Backbone Combinations
In the pursuit of identifying optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 6. The combination of Llama-3-8B as the table formatter and DeepSeek-7B as the tool maker performs the most effective pairing, having the best average execution rate and accuracy (76.8 and 68.6 respectively). This best combination aligns with our expectations given that Llama-3-8B excels in processing long tables while DeepSeek-7B, with its pre-training on code, demonstrates superior capability in tool creation.
@@
Llama-2 Llama-2 64.9/79.5 65.8/59.2 55.1/60.2 90.4/91.8 65.4/26.0 68.3/63.3
Llama-2 Llama-3 70.7/75.9 73.2/65.1 64.0/46.5 91.0/93.6 60.6/37.7 71.9/63.8
Llama-2 Codellama 70.2/76.5 73.8/74.5 64.6/56.5 94.7/88.8 71.8/34.1 75.0/66.1
Llama-2 Deepseek 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7 75.7/66.3
@@
Llama-3 Llama-3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5 74.0/65.7
Llama-3 Codellama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8 75.3/69.0
Llama-3 Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0 76.8/68.6
@@
Table 6: The TART framework with different backbone modules. The best performance is bold. The second best performance is underlined.
@@ Appendix C Tool Use on Different Backbone Models
Table 7 shows the top-10 tools that dominate the table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), consistent with our earlier findings in Section 4.1. Further illustrating this, Figure 3 (b) presents a tool categorization for the top 30 functions. We can see that table preprocessing tools constitute the highest percentage at 71.0%, followed by numerical reasoning tools at 21.8%. Together, these categories account for over 90% of tool usage, verifying our assumption that TART is better at table preprocessing and numerical reasoning.
@@
Table 7: Comparison of the top 10 functions across TART-Llama2-7b, TART-Llama3-8b, and TART-DeepSeek-7b
@@ Appendix D CoT Baseline Implementation
For a direct and fair comparison with TART, the same number of CoT samples are generated using the same IDs from the TART training dataset. These samples are generated using GPT-4, prompted with two ICL examples (detailed in Appendix F). In total, we have 9,916 training instances.
Similarly to TART, the CoT baseline was implemented across four different backbone models: Llama2-7b-hf, Llama3-8b, CodeLlama-7b-hf, and DeepSeek-Coder-7b-Instruct-V1.5. Each model was instructed to generate a step-by-step reasoning explanation followed by the final answer as per the instructions: INSTRUCTION: Given the following table, and question, generate a step-by-step reasoning explanation and the final answer.
The training process was aligned with that of TART to ensure experimental consistency. Llama-2-7b-hf, CodeLlama-7b-hf, and deepseek-coder-7b-instruct-v1.5 each requires a single GPU for approximately 12 hours, using a batch size of 4, a learning rate of 5e-5, a sequence length of 1500, gradient accumulation steps of 2, and 10 training epochs. Training Llama-3-8b requires up to 2 GPUs for around 10 hours with the same settings.
@@ Appendix E Error Analysis
To precisely categorize error types in CoT reasoning, we annotate 50 randomly selected CoT error cases. The result (Figure 7) shows that the major error type is incorrect numerical reasoning, followed by errors related to table operations. This analysis verifies the necessity for our proposed TART, which addresses these issues by integrating specialized numerical and table operation tools.
Figure 7: The error types and their distribution of CoT reasoning and TART framework.

### v3

@@ 5. SCITAB Table Fact Checking Scientific Articles Table, Claim Label (Short)
Table 5: Statistics of the seed datasets for TART training, highlighting their respective tasks, domains, and the nature of input and output data.
To construct the ToolTab, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, respectively. The detailed statistics is provided in Table 6.
@@
Table 6: Statistics in dataset ToolTab for training TART model.
@@ Appendix B Different Backbone Combinations
In the pursuit of identifying optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 7. The combination of Llama3-8b as the table formatter and DeepSeek-7b as the tool maker performs the most effective pairing, having the best average execution rate and accuracy (76.8 and 68.6 respectively). This best combination aligns with our expectations given that Llama3-8b excels in processing long tables while DeepSeek-7b, with its pre-training on code, demonstrates superior capability in tool creation.
@@
Llama2 Llama2 64.9/79.5 65.8/59.2 55.1/60.2 90.4/91.8 65.4/26.0 68.3/63.3
Llama2 Llama3 70.7/75.9 73.2/65.1 64.0/46.5 91.0/93.6 60.6/37.7 71.9/63.8
Llama2 CodeLlama 70.2/76.5 73.8/74.5 64.6/56.5 94.7/88.8 71.8/34.1 75.0/66.1
Llama2 Deepseek 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7 75.7/66.3
@@
Llama3 Llama3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5 74.0/65.7
Llama3 CodeLlama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8 75.3/69.0
Llama3 Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0 76.8/68.6
@@
Table 7: The TART framework with different backbone modules,highlighting the best (bold) and the second-best (underlined) results.
@@ Appendix C Tool Use on Different Backbone Models
Table 8 show the top 10 tools dominate the table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), consistent with our earlier findings in Section 4.1. Further illustrating this, Figure 4 (b) presents a tool categorization for the top 30 functions. Table preprocessing tools constitute the highest percentage at 71.0%, followed by numerical reasoning tools at 21.8%. Together, these categories account for over 90% of tool usage, verifying our assumption that TART is better at table preprocessing and numerical reasoning.
@@
Table 8: The top 10 functions across TART Llama2-7b, TART Llama3-8b, and TART DeepSeek-7b.
@@ Appendix D CoT Baseline Implementation
For a direct and fair comparison with TART, the same number of CoT samples are generated using the same IDs from the TART training dataset. These samples are generated using GPT-4, prompted with two in-context examples (detailed in Appendix F). In total, we obtain 9,916 training instances.
Similar to TART, the CoT baseline was implemented across four different backbone models: Llama-2-7b-hf, Llama3-8b, CodeLlama-7b-hf, and DeepSeek-Coder-7b-Instruct-V1.5. Each model was instructed to generate a step-by-step reasoning explanation followed by the final answer as per the instructions: INSTRUCTION: Given the following table, and question, generate a step-by-step reasoning explanation and the final answer.
The training process was aligned with that of TART to ensure experimental consistency. Llama-2-7b-hf, CodeLlama-7b-hf, and deepseek-coder-7b-instruct-v1.5 each requires a single GPU for approximately 12 hours, using a batch size of 4, a learning rate of 5e-5, a sequence length of 1500, gradient accumulation steps of 2, and 10 training epochs. Training Llama3-8b requires up to 2 GPUs for around 10 hours with the same settings.
Figure 8: The error types and their distributions of CoT method and our TART framework.
@@ Appendix E Error Analysis
To precisely categorize error types in CoT reasoning and TART, we annotate 50 randomly selected error cases for each method. The results (Figure 8) shows that the major error type is incorrect numerical reasoning, followed by errors related to table operations. This analysis verifies the necessity for our proposed TART, which addresses these issues by integrating specialized numerical and table operation tools.